If 𝑭 = πŸ”π’™π’›π’Š βˆ’ πŸπ’š 𝟐 𝒋 + π’šπ’›π’Œ, evaluate ∬ 𝑭 βˆ™ 𝒏𝒅𝑺 𝑺 where S is the surface of the cube bounded by 𝒙 = 𝟎, 𝒙 = 𝟏, π’š = 𝟎, π’š = 𝟏, 𝒛 = 𝟎, 𝒛 = 𝟏


1
Expert's answer
2021-06-03T17:01:14-0400

Use the Divergence Theorem


"\int\int_S\vec F\cdot\vec n\ ds=\int\int\int_E\text{div}\vec F(x, y, z)dV"

"\text{div}\vec F(x, y, z)=\nabla\cdot \vec F"

"=\dfrac{\partial}{\partial x}(6xz)+\dfrac{\partial}{\partial y}(-2y^2)+\dfrac{\partial}{\partial z}(yz)"

"=6z-4y+y=6z-3y"

Then by the Divergenve Theorem we have


"\int\int_S\vec F\cdot\vec n\ ds=\int\int\int_E\text{div}\vec F(x, y, z)dV"

"=\displaystyle\int_{0}^{1}\displaystyle\int_{0}^{1}\displaystyle\int_{0}^{1}(6z-3y)dzdydx"

"=\displaystyle\int_{0}^{1}\displaystyle\int_{0}^{1}\big[3z^2-3yz\big]\begin{matrix}
 1 \\
 0
\end{matrix}dydx"

"=3\displaystyle\int_{0}^{1}\displaystyle\int_{0}^{1}(1-y)\begin{matrix}
 1 \\
 0
\end{matrix}dydx"

"=3\displaystyle\int_{0}^{1}\big[y-\dfrac{y^2}{2}\big]\begin{matrix}
 1 \\
 0
\end{matrix}dx"

"=\dfrac{3}{2}\displaystyle\int_{0}^{1}dx"

"=\dfrac{3}{2}[x]\begin{matrix}
 1 \\
 0
\end{matrix}"

"=\dfrac{3}{2}"

"\int\int_S(6xz\vec i-2y^2\vec j+yz\vec k)\cdot\vec n\ ds=\dfrac{3}{2}"